Leo and Christian are twins. At the age of 30, Leo left for a round trip to a distance star in a spaceship with a speed of 0.95c relative to Earth. The whole trip took 20 years according to the shipboard clock. Find the ages when Leo returns

With solution


1
Expert's answer
2021-05-24T18:03:17-0400
"t=\frac{20}{1-0.95^2}=205\ years\\t_L=30+20=50\ years\\t_C=205+30=235\ years"